The formula is p=nc-e. If p=3750, n=3000, and e=900, what is c?
Answer by checkley79(3341)

p=nc-e
p=3750, n=3000, e=900
3750=3000C-900
3000C=3750+900
3000C=4650
C=4650/3000
C=1.55 ANS.